An automobile rounds a curve of radius 50.0m on a flat road at a speed of 16m/s. What centripetal acceleration is necessary to keep the car on the curve?
1
Expert's answer
2020-09-24T11:07:15-0400

Centripetal acceleration, necessary to keep the car on the curve is "a = \frac{v^2}{R}", where "v" is the speed, and "R" is the radius of the curve. Hence, "a = \frac{16^2 \frac{m^2}{s^2}}{50 m} = 5.12 \frac{m}{s^2}".